# The 2nd FutureDial Challenge: Dialog Systems with Retrieval Augmented Generation (FutureDial-RAG)

### v1

@@ 6 Challenge Rules
  • The challenge website is http://futuredial.org/. Teams should submit the registration form to FutureDialRAG@gmail.com, which will be reviewed by the organizers.
  • Teams are required to sign an Agreement for Challenge Participation and Data Usage. Data will be provided to approved teams.
  • For teams that participate in Track 1, the scores will be ranked according to the performance for Track 1. The teams can choose to participate only in Track 1.
  • For teams that participate in Track 2, they can use the baseline system provided by the organizers or use the system developed by themselves for Track 1. The ranking is based on the performance for Track 2.
  • Participants need to strictly follow the Submission Guidelines as described below. Participants are allowed to use any external (publicly available) or internal (proprietary) datasets, resources and pre-trained models,
  • The evaluation data will not be released to the teams for their own evaluation. The organizers will run the submitted systems for evaluation. The evaluation data will be shared with the eligible teams after evaluation results are announced. Only teams who strictly follow the Submission Guidelines are viewed as eligible.
  • In publishing the results, all teams will be identified as team IDs (e.g. team1, team2, etc). The organizers will verbally indicate the identities of all teams at the Workshop for communicating results. Participants may identify their own team label (e.g. team5) and report their own result, in publications or presentations, if they desire.
7 Submission Guidelines
  • Each team needs to submit a package via email to FutureDialRAG@gmail.com before the Entry Submission Deadline. The package should contain a clear README documentation for running the system over the evaluation data. The submitted system should be in one of the following two forms. In either form, the system’s processing speed should be no less than 10 tokens per second.
    • The submission package contains the system executable with the model, for example, in a Docker image. All dependencies are contained in the submission package. The organizers run the system over a server with Nvidia A100*4 hardware, evaluate, and calculate the running time over the evaluation data.
    • The system is encapsulated as a callable web service. The organizers will run the script submitted by the team, call the web service to evaluate, and calculate the running time over the evaluation data.
  • The submission should provide a System Description Document (SDD), introducing the submitted system. Teams are also encouraged to submit papers to SLT 2024. See important dates and instructions at SLT 2024 website https://2024.ieeeslt.org/.
  • Before the Entry Submission Deadline, each team can submit for multiple times for each track. The last entry from each team will be used for the evaluation.
8 Important Dates
  • April 9, 2024: Registration opening for the challenge
  • April 29, 2024: Training data release
  • June 10, 2024: Entry submission deadline
  • June 20, 2024: Evaluation results announced
  • June 20, 2024: SLT paper submission deadline
  • June 27, 2024: SLT paper update deadline
  • August 30, 2024: Notification of paper acceptance
  • December 2-5, 2024: SLT 2024 Workshop Date (in-person)
Appendix A Appendix
{CJK}
UTF8gbsn
Table 2: Chinese description for Api_query annotation.
主类 Api_query 解释
QA类 [QA] 查询FAQ手册,包含一些常用问题,如最近优惠的套餐、普遍的业务规则等。
置空类 - 根据上下文信息,客服人员无需进行额外的查询便能顺利的完成对话。
API-查询类 查询特定业务信息 查询移动当前有的业务信息,如特定的套餐、流量包等。
查询用户已办理的业务 查询用户当前已经拥有的业务,包括当前套餐、当前月租、当前流量等。
查询其他信息(例如:查询流量短信) 查询其他用于完成对话的关键信息。比如:查询历史轨迹中移动10086给用户发送的超出流量提醒短信、查询营业厅地址等
API-取消类 取消 取消用户当前拥有的某个业务
API-办理类 办理 为用户办理某个新的业务
API-验证类 验证 向用户发送验证码、密码等相关的客服验证操作
Figure 3: An example of annotated Chinese dialogs.

### v2

@@ 7 Challenge rules
Each team should submit the results for the challenge via the official email before the Entry Submission Deadline. The submission to the challenge is required to contain a system executable with all the dependencies or an encapsulated web service to be called, with a clear README documentation for running the system over the evaluation data. In either manners, the system’s processing speed should be no less than 10 tokens per second. The submission should also provide a System Description Document (SDD), introducing the submitted system. The organizers of the challenge either run the systems over a server with Nvidia A100*4 hardware, or call the web service to evaluate the submitted systems and obtain the final scores.
8 Conclusion
We present the FutureDial-RAG challenge at SLT 2024, which aims at promoting the study of RAG for dialog systems. To support the challenge, we build the MobileCS2 dataset, a real-life customer service datasets with over 3000 high-quality dialogs, containing annotations for knowledge base query and corresponding results. Upon the dataset, we define two tasks, track 1 for knowledge retrieval and track 2 for response generation, which are core research questions in dialog systems with RAG. We build baseline systems for the two tracks and design metrics to measure whether the system can perform accurate retrieval and generate informative and coherent response. The baseline results show that it is very challenging to perform well on the two tasks. We hope that the dataset and the baseline systems will be helpful to foster the community to study how to make better use of RAG for dialog systems.